OpenStudy (anonymous):

what is (3x-3) to the second power

OpenStudy (anonymous):

hi :)

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

\[(3x-3)(3x-3)=9x ^{2} -18x +9\]

OpenStudy (anonymous):

\[(3x-3)^{2} = (3x-3)(3x-3)\]

OpenStudy (anonymous):

x is multiplication

OpenStudy (anonymous):

oh okay :)

OpenStudy (anonymous):

\[(3\times-3)(3\times-3)\]

OpenStudy (anonymous):

\[3\times-3=-9\]

OpenStudy (anonymous):

\[(-9)\times(-9)\]

OpenStudy (anonymous):

\[81\]

OpenStudy (anonymous):

did you understand? :)

OpenStudy (anonymous):

so minus 10 would be 71?

OpenStudy (anonymous):

81-10=71 yes :)